OpenStudy (anonymous):

Simplify the following quotient of complex numbers into the form a + bi. -4-3i/2-i

OpenStudy (anonymous):

Let's multiply by 1 in the form of 2+i/2+i. That gives -4-3i/2-i x 2+i/2+i Using foil, on top we get: (-4-3i)x(2+i) = -8 - 6i -4i - 3isquared isquared = -1, so we get -8 - 10i +3 = -5 - 10i Now on the bottom, we foil to get: (2-i)(2+i) = 4 -2i + 2i -isquared = 4 + 1 = 5 So we have -5 - 10i/5 = -5/5 - 10i/5 = -1 - 2i